OpenStudy (lgbasallote):

a serious question this time can someone give me a trigonometric substitution problem please? i'd like to practice...and for all those people who'll answer "google" or something...i need someone to check if i did it right :P and who else is better to check than the one who gave the problem himself

OpenStudy (anonymous):

@lgbasallote nice but what do u want??

OpenStudy (lgbasallote):

....a trigonometric substitution problem....

OpenStudy (lgbasallote):

it's integral calculus...if that's what you're asking

OpenStudy (anonymous):

ok i have been studying the in-depth processes of trigonometric substitution with integrals and this problem has me frusterated. \[\int\limits_{}x^2\sqrt{(x^2-4)}dx \] The evaluation is clear (from an old Table of Integrals I found), but the derivation is not at all clear, which is what i want to know. I also tried to solve this by integration by parts, but every approach ended with an even more complicated integral, so trig substitution is probably the best choice.

OpenStudy (callisto):

Perhaps you can try these two? (well.. I haven't done that neither :P 0

OpenStudy (blockcolder):

Try finding the area of an ellipse with major axis a and minor axis b, if you don't know it yet.

OpenStudy (lgbasallote):

ill do yours first @Callisto it's easier lol..well the second one is the same so ill do the first...

OpenStudy (lgbasallote):

not that yet @blockcolder :P

OpenStudy (blockcolder):

But it's good practice! =))

OpenStudy (lgbasallote):

\(\LARGE \int \frac{1}{1+ 4x^2} dx\) let \(\tan \theta = 2x\) \(\sec^2 \theta d\theta = 2 dx\) \(\sec \theta = \sqrt{1 + 4x^2}\) \(\sec^2 \theta = 1+ 4x^2 \) substituting those.... \(\LARGE \frac{1}{2} \int \frac{\sec^2 \theta d\theta}{\sec^2 \theta} \) \(\frac{1}{2} \int d\theta\) \(\frac{1}{2} (\theta)\) since \(\tan \theta = 2x\) \(\theta = \tan^{-1} 2x\) so... \(\LARGE \frac{\tan^{-1} (2x)}{2}\)

OpenStudy (lgbasallote):

that's number 1 @Callisto

OpenStudy (anonymous):

@Rohangrr , The solution to your problem is add and subtract 4. \[(x^2 - 4) \sqrt{x^2-4}+4\sqrt{x^2-4}\] Can you do it from here or you want me to proceed??

OpenStudy (anonymous):

nope @shivam_bhalla but still tryin' i asked my teacher as well!

OpenStudy (callisto):

Yes, but you forgot to add '+C'. ( Just my opinion only. )

OpenStudy (callisto):

I meant the answer. ( you know I don't do well in integration!)

OpenStudy (lgbasallote):

oh yeah that too.. I LOST MY LATEX ON THE SECOND PROBLEM T_T

OpenStudy (lgbasallote):

to rohangrr's problem i go

OpenStudy (lgbasallote):

first step...the triangle...  from that we have \(2 \sec \theta = x\) \(2 \tan \theta = \sqrt {x^2 - 4}\) \(2 \sec \theta \tan \theta d\theta = dx\) so...trig sub... \(\int (2\sec \theta)^2 (2\tan \theta) (2 \sec \theta \tan \theta d \theta)\) \( 16 \int (\sec^2 \theta) (\tan \theta) (\sec \theta \tan \theta d\theta)\) \( 16 \int (\sec^2 \theta) (\tan ^2 \theta) (\sec \theta d\theta)\) \( 16 \int (\sec^2 \theta) (\sec ^2 \theta - 1) (\sec \theta d\theta)\) \( 16 \int (\sec^3 \theta) (\sec ^2 \theta - 1)d \theta\) \( 16 \int (\sec ^5 \theta - \sec ^3 \theta)d \theta\) that's integrable right @Rohangrr

OpenStudy (lgbasallote):

im going to redo callisto's problem now...

OpenStudy (anonymous):

Ok. I will complete it for you \[\int(x^2-4)^{3/2} + 4 * \int\sqrt{x^2- 2^2}\] From here I will do the integreal for each part separartely

OpenStudy (anonymous):

Or look at @lgbasallote solution. It is better than what I have done here on paper :)

OpenStudy (lgbasallote):

\(\int (y\sqrt {1 - y^2})dy\) let \(\cos \theta = y\) \(-\sin \theta d \theta = dy\) \(\sin \theta = \sqrt{1 - y^2}\) \(\int (\cos \theta)(\sin \theta)(-\sin \theta d\theta)\) \(-\int (\cos \theta)(\sin^2 \theta)d\theta\) \(-\int (\cos \theta)(1 - \cos^2 \theta)d\theta\) \(-\int (\cos \theta - \cos^3 \theta)d\theta\) \(-\int \cos \theta d\theta + \int \cos^3 \theta d\theta\) \(-\sin \theta + \int (1 - \sin^2 \theta)(\cos \theta) d\theta\) let \(\sin \theta =\ s\) \(\cos \theta d\theta = ds\) \(-\sin \theta + \int (1 - s^2)ds\) \(\LARGE -\sin \theta + s - \frac{s^3}{3}\) \(\large -\sin \theta + \sin \theta - \frac{\sin^3 \theta}{3}\) \(\large \frac{\sin ^3 \theta}{3}\) \(\Large \frac{(1 - y^2)^{\frac{3}{2}}}{3}\) @Callisto

OpenStudy (anonymous):

The answer to Callisto's second problem is \[\int\limits_{}^{} \ln(1+4x^2) dx\] Applying u dv rule which is \[\int\limits_{}^{} u dv = u \int\limits_{}^{}dv + v \int\limits_{}^{}du \] Here u=ln(1+4x^2) dv = dx

OpenStudy (anonymous):

\[I=\ln(1+4x^2)\int\limits_{}^{}dx + \int\limits_{}^{}x * (d/dx)(\ln(1+4x^2)\] \[I=xln(1+4x^2)+ \int\limits_{}^{}8x^2 dx / (1+4x^2) \] \[I=xln(1+4x^2) + \ln(1+4x^2) +C\] \[I=(x+1)(\ln(1+4x^2)) +C\]

OpenStudy (anonymous):

Soory It is (x-1)(ln (1+4x^2))+C I made a sign mistake in second term of udv rule . It should be negative

OpenStudy (callisto):

I did it like this: \[\int y\sqrt{1-y^2}dy \] Let y=sinθ, dy = cosθ dθ \[\int sinθ\sqrt{1-sin^2θ}cosθdθ = \int sinθcos^2θdθ = -\int cos^2θ d(cosθ)\]\[ = (-1/3) cos ^3 θ +C = [-(1-y^2)^{(3/2)}] \ /3 +C\]

OpenStudy (callisto):

Sorry, I type slow :S

OpenStudy (lgbasallote):

lol i guess that was faster =))

OpenStudy (lgbasallote):

i meant the method

OpenStudy (callisto):

You forgot a '-' sign for your answer :S

OpenStudy (anonymous):

@Callisto , nobody can defeat me in typing slowly :P BTW, your method was however faster :) and yes you forgota a negative sign :P

OpenStudy (callisto):

@lgbasallote Check your last 3 steps

OpenStudy (lgbasallote):

oh yeah lol...it's just typing latex so long makes you skip those :p =)))

OpenStudy (callisto):

@shivam_bhalla Well... Now you know who can :)

OpenStudy (anonymous):

Haha :P I think we give each other a close competition in typing you know :P BTW, Do you want to solve Qn 5 part b) ?? @Callisto

OpenStudy (lgbasallote):

\(\Huge \color{forestgreen}{\mathtt{modesty}}\) \(\Huge \color{skyblue}{\mathcal{WARS!!!}}\)

OpenStudy (callisto):

@shivam_bhalla That's easy :S

OpenStudy (anonymous):

LOLOLOL.............

OpenStudy (lgbasallote):

wow...and you said you werent good in integration :P

OpenStudy (callisto):

@lgbasallote Then, I lose :) Mine is not modesty, it is honesty :)

OpenStudy (callisto):

@lgbasallote I'm not... Certainly not!

OpenStudy (callisto):

(not good at doing integration problems!)

OpenStudy (anonymous):

@lgbasallote you said in the beginning " this time <silence> " LOL. We are violating these conditions to the core :P

OpenStudy (lgbasallote):

ah we are opposite..im better in these problems lol @shivam_bhalla lol it was a joke..serious -> silence haha

OpenStudy (anonymous):

Ok. Any more tough problems?? Or I shall give tough problems??

OpenStudy (lgbasallote):

i am assuming that question is not directed for me as i have not given any problem so i doubt i should give MORE

OpenStudy (callisto):

Do you really want to try something difficult?

OpenStudy (anonymous):

Well this is for everyone :) Evaluate the following expression : \[\int\limits_{}^{}dx/(\sin^4 x+ \cos^4 x)\]

OpenStudy (anonymous):

@Callisto , Yes

OpenStudy (lgbasallote):

@shivam_bhalla that is trig sub right?

OpenStudy (anonymous):

Well, not exactly. No trig sub

OpenStudy (anonymous):

But yes you have to use trig identities to solve this :)

OpenStudy (callisto):

I'm not sure.. I seem to have done something like that this morning :|

OpenStudy (callisto):

Should I apply double angle formula here?

OpenStudy (anonymous):

No. I will give you a tip. Multiply and divide by sec^4 (x) and in numerator split sec^4 (x) = sec^2 (x)* sec^2 (x) =(1+tan^2 (x))*(sec^2 (x)) Now try to solve. @Callisto

OpenStudy (anonymous):

tip 2 = Multiply the sec^4x in the denominator and try to bring it in only tan form i.e all the terms must be tan and not sec. Then take tan x = t and try to solve :)

OpenStudy (callisto):

So many tips.... :S

OpenStudy (callisto):

It's going to be something ugly :S

OpenStudy (anonymous):

It is just the trailer. :P The film is still remaining :P

OpenStudy (callisto):

I don't have a good feeling... (Shivering)

OpenStudy (anonymous):

Be positive. Take it as a challenge to solve this :) Say it is possible. Definitely you will arrive at the solution